a box contains five ball. the balls are numbered 1,2,3,4 and 5. a person randomly draws to balls from the box. let X be the sum of the numbers kf the ball drawn. what is the sum of all probabilities, P(X=x)?


1
Expert's answer
2022-02-21T16:40:34-0500

There are "\dbinom{5}{2}=10" outcomes.


"\def\arraystretch{1.5}
 \begin{array}{c:c}
 Sample\ point & X \\
 \hdashline
 1,2 & 3 \\
 \hdashline
1,3 & 4 \\
 \hdashline
1,4 & 5 \\
 \hdashline
1,5 & 6 \\
 \hdashline
2,3 & 5 \\
 \hdashline
2,4 & 6 \\
 \hdashline
2,5 & 7 \\
 \hdashline
 3,4 & 7 \\
 \hdashline
 3,5 & 8 \\
 \hdashline
 4,5 & 9 \\
 \hdashline
\end{array}""\def\arraystretch{1.5}
 \begin{array}{c:c}
 x & 3 & 4 & 5 & 6 & 7 & 8 & 9 \\ \hline
 p(x) & 0.1& 0.1 & 0.2 & 0.2 & 0.2 & 0.1 & 0.1 \\
\end{array}"

Check


"\sum_ip(x_i)=0.1+0.1+0.2+0.2+0.2+0.1+0.1=1"

"mean=E(X)=\sum _ix_ip(x_i)=3(0.1)+4(0.1)"

"+5(0.2)+6(0.2)+7(0.2)+8(0.1)+9(0.1)=6"